OpenStudy (anonymous):

find the following... assume that the variables can represent any real numbers... (a+7) 2 − − − − − − − √

OpenStudy (anonymous):

can you use the draw tool and put the question there ?

OpenStudy (anonymous):

\[\sqrt{(a+7)^2}\]ya

OpenStudy (anonymous):

thats wat it looks like

OpenStudy (anonymous):

\[\sqrt{(a+7)^2}=\left| a+7 \right|\]

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

can u help me with another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\sqrt[3]{-125x^3?}\]

OpenStudy (anonymous):

it looks like that but no questions mark at the end it says simplify

OpenStudy (anonymous):

\[\sqrt[3]{-125x^3}=\sqrt[3]{-1}*\sqrt[3]{5^3}*\sqrt[3]{x^3}=-5x\]

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[\sqrt[4]{2401}\]

OpenStudy (anonymous):

\[\sqrt[4]{2401} = \sqrt{49} = 7\]